OpenStudy (anonymous):

What is the slope of the line that passes through the points (–2, 5) and (1, 11)? A. B. C. D.

OpenStudy (anonymous):

your slope is m=2. in this particular instance, this is where your formula knowledge comes in handy. utilize what you know from the concepts you learned from solving and understanding slopes. In order to solve this problem, you must use the point-slope formula: m = (y2 – y1)/(x2 – x1) those numbers are supposed to be in superscript. (11 - 5)/(1 -(-2)) = 6/3 = 2

OpenStudy (anonymous):

Thanks! :3